On July 1, 2018, Gupta Corporation bought 25% of the outstanding common stock of VB Company for $160 million cash. At the date of acquisition of the stock, VB’s net assets had a total fair value of $540 million and a book value of $280 million. Of the $260 million difference, $58 million was attributable to the appreciated value of inventory that was sold during the last half of 2018, $156 million was attributable to buildings that had a remaining depreciable life of 15 years, and $46 million related to equipment that had a remaining depreciable life of 5 years. Between July 1, 2018, and December 31, 2018, VB earned net income of $60 million and declared and paid cash dividends of $52 million.

Required:
1. Prepare all appropriate journal entries related to the investment during 2018, assuming Gupta accounts for this investment by the equity method.
2. Determine the amounts to be reported by Gupta.
Answer #1

Answer:

Part 1
Date Account Titles and Explanation Debit Credit W.N
01-07-2018 Investment in Company V Shares 160,000,000
          Cash 160,000,000
To record purchase of investment
31-12-2018 Investment in Company LC Shares 15,000,000 =60000000*25%
          Investment Revenue 15,000,000
To record shares of net income as revenue from Investment
31-12-2018 Cash 13,000,000 = 52000000*25%
         Investment in Company LC Shares 13,000,000
To record receipt of dividends
31-12-2018 Investment Revenue 16,950,000 W.N 3
          Investment in Company V Shares 16,950,000
To record amortization adjustment
W.N 3
Details Amount Amount Calculation
Share of investor in fair value of investee’s net assets 135,000,000 =(540,000,000*25%)
Share of investor in book value of investee’s net assets 70,000,000 =(280,000,000*25%)
Difference of depreciable net assets 65,000,000
Amount attributed to inventory 14,500,000 =(58,000,000*25%)
Amount attributed t building 39,000,000 =(156,000,000*25%)
Deprecation of Building 1,300,000 =39,000,000/15*6/12
Amount attributed to equipment 11,500,000 =(46,000,000*25%)
Depreciation of Equipment 1,150,000 =11,500,000/5*6/12
Total Amortization 16,950,000 =14,500,000+1,300,000+1,150,000
Part 2
Investment in Company V Shares
Details Debit Details Credit
Cost 160,000,000 Dividends 13,000,000
Share of Income 15,000,000 Depreciation 16,950,000
Total 175,000,000 Total 29,950,000
Balance 145,050,000
Investment Revenue
Details Debit Details Credit
Depreciation 16,950,000 Share of Income 15,000,000
Total 16,950,000 Total 15,000,000
Balance 1,950,000